# Rendering-Oriented 3D Point Cloud Attribute Compression using Sparse Tensor-based Transformer

### v2

@@ IV-B Compression Performance Comparison
Class Point Cloud TMC13 v14 [43] TMC13 v23 [43] SparsePCAC [21] ScalablePCAC [22]
Y YUV MS-SSIM Y YUV MS-SSIM Y YUV MS-SSIM Y YUV MS-SSIM
@@
soldier -27.90 -27.37 -32.67 -20.39 -19.60 -32.67 -28.21 -28.44 -28.03 -11.35 -10.92 -16.17
Average -22.66 -16.60 -11.15 -17.06 -9.62 -4.13 -24.65 -26.51 -28.03 -1.74 -2.46 -10.69
Owlii basketball player -33.62 -32.51 -41.09 -25.81 -24.17 -33.34 -35.17 -35.21 -31.36 -12.05 -10.71 -20.15
dancer -31.44 -30.51 -36.66 -24.85 -23.54 -30.54 -29.04 -29.43 -31.27 -9.21 -8.50 -15.45
@@
Average -29.85 -29.04 -37.86 -22.82 -21.46 -31.12 -31.18 -31.56 -28.31 -10.37 -9.61 -18.32
@@
Furthermore, our method was also compared with SparsePCAC and ScalablePCAC, two other learning-based models for point cloud attribute compression. To ensure a fair comparison, SparsePCAC was trained according to the training settings of this study. Due to the source code of ScalablePCAC not being open and the use of an identical training dataset in our study, we asked the author for the reconstructed point clouds at different ratepoints to compare. Both methods are tested under the same conditions. As shown in Table. I, our method achieves an average BD-BR reduction of 24.65% (26.51%) and 31.18% (31.56%) over SparsePCAC in the two test datasets. Compared to ScalablePCAC, our method also achieves a BD-BR reduction of 1.74% (2.46%) and 10.37% (9.61%), respectively.
In addition, we have employed the MS-SSIM metric for a more comprehensive assessment of performance. Our method surpasses the G-PCC TMC13v14, TMC13v23, SparsePCAC and ScalablePCAC by 11.15%, 4.13%, 28.03% and 10.69% BD-BR in 8iVFB, respectively. In the Owlii dataset, the respective improvements are 37.86%, 31.18%, 28.31%, and 18.32%.
@@
Class Point Cloud TMC13 v14 [43] TMC13 v23 [43] SparsePCAC [21] ScalablePCAC [22]
$128^{2}$ $256^{2}$ $512^{2}$ $128^{2}$ $256^{2}$ $512^{2}$ $128^{2}$ $256^{2}$ $512^{2}$ $128^{2}$ $256^{2}$ $512^{2}$
@@
soldier -35.09 -35.89 -33.43 -28.45 -27.89 -25.47 -28.83 -27.81 -27.76 -11.45 -10.92 -10.58
Average -22.81 -23.38 -21.23 -16.74 -16.00 -13.60 -26.36 -26.37 -26.19 -3.63 -3.25 -3.14
Owlii basketball player -31.86 -31.95 -31.39 -21.40 -22.51 -22.77 -34.93 -33.25 -33.41 -9.23 -9.97 -10.19
dancer -29.42 -29.40 -30.35 -18.79 -19.99 -22.95 -29.38 -28.45 -28.86 -7.69 -7.02 -8.34
@@
Average -25.25 -26.76 -29.02 -17.13 -19.30 -21.28 -28.27 -29.04 -30.46 -6.75 -7.22 -9.70
To accommodate the varying resolution requirements of display devices, we also conducted tests to evaluate the generalization of RO-PCAC across different resolutions. Experimental results of image resolutions of 128$\times$128, 256$\times$256 and 512$\times$512 are represented in Table. II. We can see that the advantages of RO-PCAC method are still significant compared to other methods, which demonstrates that the neural network has learned suitable parameters to adapt to the rendering mechanism.
@@ IV-C Ablation Study
In this section, we present an ablation study to evaluate the contributions of SP-Trans and Rendering Module in our proposed model. The ablation study is conducted by incrementally removing components from the full model and measuring the resulting performance degradation. The outcomes of the ablation study are shown in Fig. 14 and Table. III, which provide a comparative analysis of the model’s performance without SP-Trans and Rendering Module (i.e., w/o SP-Trans and w/o RM). Specially, when the Rendering Module is removed, point-to-point loss is used to train the corresponding model. The exclusion of SP-Trans, Rendering Module and both of them resulted in a decrement in BD-BR by 9.86%, 7.38% and 17.02% in Y component in 8iVFB, respectively. Similarly, on the Owlii dataset, the absence of these components led to decrements of 10.94%, 8.83%, and 20.20% in the BD-BR for the Y component, respectively, which demonstrates the effectiveness of SP-Trans and the Rendering Module.
@@
Table III: Ablation Studies on Modular Components in RO-PCAC
@@
8iVFB longdress 2.40 2.21 4.69
redandblack 6.11 6.38 12.30
loot 12.78 11.22 24.67
soldier 18.15 9.65 26.43
Average 9.86 7.38 17.02
Owlii basketball player 12.56 9.27 22.19
dancer 8.85 8.07 17.58
exercise 17.46 11.28 29.66
model 4.89 6.69 11.37
Average 10.94 8.83 20.20
@@ IV-D Compression Complexity
Table IV illustrates the detailed time consumption in the encoder and decoder. For learning-based methods, ‘Y and Z analysis’ represents the time of generating the latent and hyperprior features in the encoder and ‘Y and Z synthesis’ represents the time of reconstructing point clouds through latent and hyperprior features in the decoder. ‘Y and Z compressing’ and ‘Y and Z decompressing’ represent the time of compressing and decompressing the latent and hyperprior features by the arithmetic encoder and arithmetic decoder, respectively. ‘B.’ and ‘E.’ represent the base layer and enhancement layer in ScalablePCAC, respectively. We can observe that RO-PCAC has a comparable encoding/decoding time consumption with G-PCC, which are 4.42s/4.79s and 3.92s/3.68s, respectively. However, the encoding/decoding times of SparsePCAC and ScalablePCAC are much longer than RO-PCAC, which are 20.79s/201.52s and 53.68s/351.54s, respectively. This is due to the entropy model they adopted needing to take previous decoded latent features as input to generate the entropy model parameters of the current features to be encoded. In contrast, RO-PCAC directly outputs the entropy parameters of all latent features and encode them simultaneously, having a huge time saving. In addition, in the learning included methods, SparsePCAC has the shortest time of feature analysis and synthesis, which are 0.17s and 0.44s, respectively. Compared with SparsePCAC, RO-PCAC only has time addition of 0.1s and 0.09s, respectively. This is attributed to the efficient neighbor searching in SP-Trans, which has the same time complexity as that in sparse convolution.
Table IV: Runtime (seconds) analysis
@@
We have presented RO-PCAC, a deep learning framework for rendering-oriented point cloud compression. By integrating the differentiable rendering into the compression process, RO-PCAC balances compression efficiency with the quality of rendered images. The introduction of SP-Trans, leveraging sparse tensor representations, efficiently enhances feature analysis and synthesis by capturing the complex interdependencies in regions of varying density. Our experimental results indicate that RO-PCAC achieves the state-of-the-art performance, demonstrating improvements in compression efficiency compared to the MPEG G-PCC standard and other contemporary learning-based and hybrid methods.

### v3

@@ IV-A3 Baselines
  • TSC-PCAC [45]. This is an end-to-end voxel transformer-based point cloud compression method.
@@ IV-B Compression Performance Comparison
Class Point Cloud TMC13 v14 [43] TMC13 v23 [43] SparsePCAC [21] TSC-PCAC [45] ScalablePCAC [22]
Y YUV MS-SSIM Y YUV MS-SSIM Y YUV MS-SSIM Y YUV MS-SSIM Y YUV MS-SSIM
@@
soldier -27.90 -27.37 -32.67 -20.39 -19.60 -32.67 -28.21 -28.44 -28.03 -10.06 -12.56 4.95 -11.35 -10.92 -16.17
Average -22.66 -16.60 -11.15 -17.06 -9.62 -4.13 -24.65 -26.51 -28.03 -15.07 -15.40 -14.75 -1.74 -2.46 -10.69
Owlii basketball -33.62 -32.51 -41.09 -25.81 -24.17 -33.34 -35.17 -35.21 -31.36 -12.12 -11.35 -1.05 -12.05 -10.71 -20.15
dancer -31.44 -30.51 -36.66 -24.85 -23.54 -30.54 -29.04 -29.43 -31.27 -9.26 -9.44 -2.84 -9.21 -8.50 -15.45
@@
Average -29.85 -29.04 -37.86 -22.82 -21.46 -31.12 -31.18 -31.56 -28.31 -10.69 -10.40 -1.95 -10.37 -9.61 -18.32
‘–’ indicates the test point cloud is selected into training dataset.
@@
Furthermore, our method was also compared with TSC-PCAC, SparsePCAC and ScalablePCAC, three other learning-based models for point cloud attribute compression. To ensure a fair comparison, SparsePCAC was trained according to the training settings of this study. In the same way, we retrained TSC-PCAC using our training conditions, but the performance was not as good as the trained models provided by the authors. Therefore, we used their trained models to report their performance in our paper. Due to the source code of ScalablePCAC not being open and the use of an identical training dataset in our study, we asked the author for the reconstructed point clouds at different ratepoints to compare. All methods are tested under the same conditions. As shown in Table. I, our method achieves an average BD-BR reduction of 24.65% (26.51%) and 31.18% (31.56%) over SparsePCAC in the two test datasets. Compared to TSC-PCAC, our method achieves a BD-BR reduction of 15.07% (15.40%) and 10.69% (10.40%), respectively. Compared to ScalablePCAC, our method also achieves a BD-BR reduction of 1.74% (2.46%) and 10.37% (9.61%), respectively.
In addition, we have employed the MS-SSIM metric for a more comprehensive assessment of performance. Our method surpasses the G-PCC TMC13v14, TMC13v23, SparsePCAC, TSC-PCAC and ScalablePCAC by 11.15%, 4.13%, 28.03%, 14.75% and 10.69% BD-BR in 8iVFB, respectively. In the Owlii dataset, the respective improvements are 37.86%, 31.18%, 28.31%, 1.95% and 18.32%.
@@
Class Point Cloud TMC13 v14 [43] TMC13 v23 [43] SparsePCAC [21] TSC-PCAC [45] ScalablePCAC [22]
$128^{2}$ $256^{2}$ $512^{2}$ $128^{2}$ $256^{2}$ $512^{2}$ $128^{2}$ $256^{2}$ $512^{2}$ $128^{2}$ $256^{2}$ $512^{2}$ $128^{2}$ $256^{2}$ $512^{2}$
@@
soldier -35.09 -35.89 -33.43 -28.45 -27.89 -25.47 -28.83 -27.81 -27.76 -15.18 -13.68 -13.06 -11.45 -10.92 -10.58
Average -22.81 -23.38 -21.23 -16.74 -16.00 -13.60 -26.36 -26.37 -26.19 -16.60 -14.95 -14.89 -3.63 -3.25 -3.14
Owlii basketball -31.86 -31.95 -31.39 -21.40 -22.51 -22.77 -34.93 -33.25 -33.41 -7.81 -11.29 -9.90 -9.23 -9.97 -10.19
dancer -29.42 -29.40 -30.35 -18.79 -19.99 -22.95 -29.38 -28.45 -28.86 -7.28 -6.85 -8.31 -7.69 -7.02 -8.34
@@
Average -25.25 -26.76 -29.02 -17.13 -19.30 -21.28 -28.27 -29.04 -30.46 -7.55 -9.07 -9.11 -6.75 -7.22 -9.70
‘–’ indicates the test point cloud is selected into training dataset.
To accommodate the varying resolution requirements of display devices, we also conducted tests to evaluate the generalization of RO-PCAC across different resolutions. Experimental results of image resolutions of 128$\times$128, 256$\times$256 and 512$\times$512 are represented in Table. II. We can see that the advantages of RO-PCAC method are still significant compared to other methods, which demonstrates that the neural network has learned suitable parameters to adapt to the rendering mechanism. We have also observed that although RO-PCAC outperforms all other baseline methods, the gains in BD-BR is reduced for lower-resolution multi-view images compared to 1024x1024 resolution. We attribute this reduction to the resolution discrepancy between training and testing images.
In order to provide a more holistic evaluation of the proposed method, we implement an evaluation of using point-to-point Y PSNR as a 3D metric. As shown in Table III, compared with the results of image Y PSNR, the gains of all test point cloud decrease. For the 8iVFB dataset, RO-PCAC exhibits a performance that is slightly inferior to the state-of-the-art method for reconstructing the original point cloud (i.e., ScalablePCAC) by 1.17%. These outcomes align with our research emphasis, as we aim to optimize for rendered image quality, which represents the ultimate output as perceived by the end-user in point cloud visualization scenarios.
Table III: BD-BR (Y components) comparison with TMC13 v23, SparsePCAC and ScalablePCAC in point-to-point PSNR
Class Point Cloud TMC13 v23 SparsePCAC TSC-PCAC ScalablePCAC
Y Y Y Y
8iVFB longdress -13.04 -8.36 9.10
redandblack -16.12 -18.58 -13.99 1.46
loot -3.94 -33.76 2.33
soldier -20.17 -25.52 -9.76 -8.23
Average -13.32 -21.55 -11.88 1.17
Owlii basketball -25.31 -33.06 -13.28 -8.82
dancer -26.14 -26.76 -9.71 -7.05
exercise -18.69 -34.58 -3.81
model -19.64 -21.40 -9.02
Average -22.45 -28.95 -11.50 -7.18
‘–’ indicates the test point cloud is selected into training dataset.
Different distortion measures in the loss function can be used in our framework. Following the convention of image and video compression [69, 70, 71, 72], which uses Multi-Scale Structural Similarity (MS-SSIM) as loss function to enhance the subjective quality of images, we also adopted the MS-SSIM as the loss function to train the model. Specifically, the distortion term is defined by $1-\text{MS}\text{-}\text{SSIM}(I,\hat{I})$. The results shown in Table IV indicate that when the model was optimized using MS-SSIM, there were increases in performance for all test point clouds compared to optimization with MSE. This indicates that the MS-SSIM optimized model improves the structural similarity of the rendered images of point clouds.
Table IV: BD-BR (MS-SSIM) comparison with TMC13 v23, SparsePCAC and ScalablePCAC using MS-SSIM optimized RO-PCAC
Class Point Cloud TMC13 v23 SparsePCAC TSC-PCAC ScalablePCAC
MS-SSIM MS-SSIM MS-SSIM MS-SSIM
8iVFB longdress -11.36 -16.21 -10.40
redandblack 10.51 -42.18 -39.29 -13.96
loot -13.52 -43.12 -9.31
soldier -38.17 -31.17 -1.76 -18.42
Average -13.14 -33.17 -20.53 -13.02
Owlii basketball -36.21 -34.01 -3.28 -24.12
dancer -33.11 -36.58 -5.13 -17.09
exercise -29.14 -34.29 -23.31
model -36.64 -21.12 -20.62
Average -33.78 -31.50 -4.21 -21.29
‘–’ indicates the test point cloud is selected into training dataset.
Except for the human-based models, RO-PCAC can also compress indoor scene point cloud such as ScanNet [66]. This experiment reports the compression performance over G-PCC TMC13 v23. To avoid presenting a simplistic average performance of the 100 test point clouds, we report the average performance at increments of 25, specifically for the first 25, 50, 75, and all 100 point clouds, to demonstrate the robustness of our method across various indoor scenes. The results presented in Table V demonstrate that RO-PCAC outperforms G-PCC TMC13 v23 across all frame sets.
Table V: BD-BR (Y component, YUV component and MS-SSIM) comparison with TMC13 v23 of ScanNet under different frame sets
Metrics Frames
25 50 75 100
Y -10.48 -9.60 -7.87 -8.61
YUV -7.03 -6.03 -4.12 -5.19
MS-SSIM -10.45 -11.36 -9.12 -10.19
@@ IV-C Ablation Study
IV-C1 Ablation Study of SP-Trans and Rendering Module
In this section, we present an ablation study to evaluate the contributions of SP-Trans and Rendering Module in our proposed model. The ablation study is conducted by replacing or removing components from the full model and measuring the resulting performance degradation. The outcomes of the ablation study are shown in Fig. 14 and Table. VI, which provide a comparative analysis of the model’s performance through replacing SP-Trans with SConv and removing Rendering Module (i.e., w/o SP-Trans and w/o RM). Specially, when the Rendering Module is removed, point-to-point loss is used to train the corresponding model. The exclusion of SP-Trans, Rendering Module and both of them resulted in a decrement in BD-BR by 6.66%, 7.38% and 13.96% in Y component in 8iVFB, respectively. Similarly, on the Owlii dataset, the absence of these components led to decrements of 8.32%, 8.83%, and 16.94% in the BD-BR for the Y component, respectively, which demonstrates the effectiveness of SP-Trans and the Rendering Module.
@@ IV-C1 Ablation Study of SP-Trans and Rendering Module
Table VI: Ablation Studies on Modular Components in RO-PCAC
@@
8iVFB longdress 1.59 2.21 3.54
redandblack 3.71 6.38 10.20
loot 8.18 11.22 20.24
soldier 13.23 9.65 21.92
Average 6.66 7.38 13.96
Owlii basketball 10.11 9.27 19.14
dancer 6.15 8.07 14.66
exercise 13.79 11.28 24.19
model 3.21 6.69 9.76
Average 8.32 8.83 16.94
IV-C2 Ablation Study of Variations in SP-Trans
In this section, we present an ablation study to evaluate the contributions of variations in SP-Trans, compared with the typical point transformer [25]. The ablation study is conducted by individually replacing the variations in SP-Trans with the corresponding components in [25] and measuring the resulting performance degradation. The outcomes of the ablation study are shown in Table. VII, which provide a comparative analysis of the model’s performance through replacing cosine with softmax similarity measurement and replacing neighbor searching in 3D window with K nearest neighbors (K=16 as suggested by [25]) (i.e., w/ softmax and w/ KNN). Specially, in order to demonstrate the efficient neighbor searching in 3D window of SP-Trans, we test the added runtime by using KNN. The replacement with softmax and KNN resulted in a decrement in BD-BR by 5.63% and 3.65% in Y component in 8iVFB, respectively. Similarly, on the Owlii dataset, the absence of these components led to decrements of 4.80% and 3.78% in the BD-BR for the Y component, respectively. In addition, the total runtime of encoder and decoder by using KNN added 0.41s (43%) and 0.31s (33%), respectively.
Table VII: Ablation Studies on Variations in SP-Trans
Class Point Cloud w/ softmax w/ KNN
Y Y Added Runtime
8iVFB longdress 3.44 4.34 0.36s (38%)
redandblack 5.26 6.67 0.58s (43%)
loot 4.26 3.21 0.27s (45%)
soldier 5.17 3.65 0.43s (47%)
Average 5.63 4.47 0.41s (43%)
Owlii basketball 2.46 3.26 0.39s (41%)
dancer 5.31 3.37 0.28s (22%)
exercise 5.16 6.28 0.21s (33%)
model 6.26 2.21 0.37s (46%)
Average 4.80 3.78 0.31 (33%)
The relative added percentage is indicated in parentheses.
IV-C3 Ablation Study of Number of Projections
In this section, we present an ablation study to evaluate the influence of the number of projection images on the performance. This study involves varying the number of projections used during the training and testing phases and presenting corresponding compression performances in dataset 8iVFB. For training projection, the numbers are set to 2, 4, 6, 8 and 14 respectively. For the 2 projections, the elevations is $0^{\circ}$ and azimuths are $0^{\circ}$ and $180^{\circ}$. For 4, 6, 8 and 14 projections, 2 views are with the elevations of $90^{\circ}$ and $270^{\circ}$ and azimuths of $0^{\circ}$, and the other views are with the elevations of $0^{\circ}$ and azimuths ranging from $0^{\circ}$ to $360^{\circ}$ with intervals of $180^{\circ}$, $90^{\circ}$, $60^{\circ}$ and $30^{\circ}$, respectively. For testing projection, the numbers are set to 2, 4, 6 and 12 respectively. In order to avoid overlapping perspectives adopted in training, the elevations is $0^{\circ}$ and azimuths ranges from $15^{\circ}$ to $360^{\circ}$ with intervals of $180^{\circ}$, $90^{\circ}$, $60^{\circ}$ and $30^{\circ}$, respectively. The outcomes of the ablation study are shown in Table. VIII, it is evident that as the number of training projections increases, the BD-BR values become more negative. However, when the number of training projections is greater than 4, the gains tend to decrease. This can be observed in the table where the BD-BR values approach -17.20% for training projections of 6, 8, and 12, with only marginal differences between these values. This suggests that beyond a certain point, which in this case is around 8 training projections, the additional training data does not significantly contribute to further performance improvements. This trend is consistent across all testing projection counts. Moreover, for a well-trained model, as the number of testing projections increases from 2 to 12, the performance remains stable with only slight changes, which indicates that the model is robust to testing projection.
The plateauing and stabilizing of gains can be attributed to the strong supervision in training. With the convention of the geometry losslessly encoded, the geometry of the decoded point clouds are the same as those of original point clouds. Furthermore, the pixels between the images respectively rendered from the original and decoded point cloud have accurate one-to-one correspondences. Based on these strong supervisions, the gradients calculated from MSE of images are allowed to be propagated from pixel values backward to latent features of exact points, making the network is easy to converge and extract effective features. Therefore, the number of projection images for training does not need too many, as long as the images can basically cover the point cloud. If the number of projection images for training is too small to cover the point cloud, features of partial points will not be supervised, and thus the trained network generates quality relatively lowered images. If more projection images are used, there also is no significant improvement in performance.
Table VIII: BD-BR (Y component) comparison with TMC13 v23 of 8iVFB under different numbers of projections
Training projections
Testing projections 2 4 6 8 14
2 -8.28 -13.19 -15.90 -17.10 -17.20
4 -7.78 -13.13 -15.92 -17.01 -17.14
6 -8.13 -12.59 -15.77 -17.13 -17.14
12 -7.97 -13.01 -15.89 -17.11 -17.17
@@ IV-D Compression Complexity
Table IX illustrates the detailed time consumption in the encoder and decoder. For learning-based methods, ‘Y and Z analysis’ represents the time of generating the latent and hyperprior features in the encoder and ‘Y and Z synthesis’ represents the time of reconstructing point clouds through latent and hyperprior features in the decoder. ‘Y and Z compressing’ and ‘Y and Z decompressing’ represent the time of compressing and decompressing the latent and hyperprior features by the arithmetic encoder and arithmetic decoder, respectively. ‘B.’ and ‘E.’ represent the base layer and enhancement layer in ScalablePCAC, respectively. We can observe that RO-PCAC has a comparable encoding/decoding time consumption with G-PCC, which are 4.42s/4.79s and 3.92s/3.68s, respectively. However, the encoding/decoding times of SparsePCAC and ScalablePCAC are much longer than RO-PCAC, which are 20.79s/201.52s and 53.68s/351.54s, respectively. This is due to the entropy model they adopted needing to take previous decoded latent features as input to generate the entropy model parameters of the current features to be encoded. In contrast, RO-PCAC directly outputs the entropy parameters of all latent features and encode them simultaneously, having a huge time saving. In addition, in the learning included methods, SparsePCAC has the shortest time of feature analysis and synthesis, which are 0.17s and 0.44s, respectively. Compared with SparsePCAC, RO-PCAC only has time addition of 0.1s and 0.09s, respectively. This is attributed to the efficient neighbor searching in SP-Trans, which has the same time complexity as that in sparse convolution.
Table IX: Runtime (seconds) analysis
@@
TSC-PCAC Y and Z analysis 0.57 -
Y and Z synthesis - 0.84
Y and Z compressing 5.62 -
Y and Z decompressing - 6.08
total 5.79 6.52
@@
We have presented RO-PCAC, a deep learning framework for rendering-oriented point cloud compression. By integrating the differentiable rendering into the compression process, RO-PCAC balances compression efficiency with the quality of rendered images. The introduction of SP-Trans, leveraging sparse tensor representations, efficiently enhances feature analysis and synthesis by capturing the complex interdependencies in regions of varying density. Our experimental results indicate that RO-PCAC achieves the state-of-the-art performance, demonstrating improvements in compression efficiency compared to the MPEG G-PCC standard and other contemporary learning-based and hybrid methods. However, it is noted that the performance gain over ScalablePCAC is relatively small, less than 10%. This motivates us to further explore more strategies to enhance RO-PCAC to achieve more gains in the future. Incorporating resolution as a variable in the framework to enhance performance for low-resolution images is also a valuable direction for future work, which would be beneficial for bandwidth limited scenarios where transmitting low-resolution images may be necessary. Moreover, the fixed weights in alpha-blending (i.e., Eq. (10)) in the rendering module may limit the model’s flexibility. Investigating adaptive weight strategies could be a valuable research to improve the rendering pipeline. This could include learning the weights dynamically based on the data or integrating attention mechanisms to allocate weights in a manner that better captures the salient features of the point cloud data.